Find two consecutive integers such that the sum of 2 times the first integer and 7 times second integer is 88.







Answer by checkley71(8403)

2x+7(x+1)=88 or 2x+7x+7=88 or 9x=81 or x=9
prood 2*9+7(10)=88 or 18+70=88 or 88=88